# Matter-of-Fact: A Benchmark for Verifying the Feasibility of Literature-Supported Claims in Materials Science

### v1

@@ Abstract
Contemporary approaches to assisted scientific discovery use language models to automatically generate large numbers of potential hypothesis to test, while also automatically generating code-based experiments to test those hypotheses. While hypotheses can be comparatively inexpensive to generate, automated experiments can be costly, particularly when run at scale (i.e. thousands of experiments). Developing the capacity to filter hypotheses based on their feasibility would allow discovery systems to run at scale, while increasing their likelihood of making significant discoveries. In this work we introduce Matter-of-Fact, a challenge dataset for determining the feasibility of hypotheses framed as claims. Matter-of-Fact includes 8.4k claims extracted from scientific articles spanning four high-impact contemporary materials science topics, including superconductors, semiconductors, batteries, and aerospace materials, while including qualitative and quantitative claims from theoretical, experimental, and code/simulation results. We show that strong baselines that include retrieval augmented generation over scientific literature and code generation fail to exceed 72% performance on this task (chance performance is 50%), while domain-expert verification suggests nearly all are solvable – highlighting both the difficulty of this task for current models, and the potential to accelerate scientific discovery by making near-term progress. [^1] Benchmark and models: https://github.com/cognitiveailab/matter-of-fact
@@ 1 Introduction
Generating data to test feasibility assessment is challenging, as (by nature) the experimental results of proposed hypotheses are as yet unknown, which makes gold labels for determining whether a hypothesis is feasible or infeasible effectively unavailable. To address this challenge, we operationalize feasibility assessment as a temporally-filtered claim verification task. As with conventional claim verification tasks (e.g. Thorne et al., 2018), we generate a corpus of claims extracted from recent scientific literature – however, in addition, each claim is paired with a “knowledge cut-off date”, which is the date that the paper the claim was generated from was first authored. When performing the feasibility assessment task, models are allowed to use any information available before the source paper was authored to assess feasibility, essentially rewinding into the past to predict future results. In this way, models are provided with knowledge up to (for example) 2023, and must use that knowledge (through a combination of literature search, small-scale code-based experimentation, world modeling, or other methods) to predict whether genuine results (and artificially-generated infeasible results) from 2024 are feasible or infeasible.
@@
MatterOfFact (This work) Mat. Sci 8.4k Paper full-text Extract positive, generate infeasible
@@ 2 Related Work
The scientific claim verification task requires a model to determine whether a claim (typically extracted from a scientific paper) is true or false, either by leveraging its pretrained scientific knowledge or retrieving evidence from a corpus, with a selection of scientific claim verification benchmarks shown in Table 1. SciFact Wadden et al. (2020) contains 1.4K biomedical-domain claims generated by showing citances (sentences that cite a paper and describe its contribution) to human annotators, who were then asked to generate associated claims. Where SciFact pairs claims with a set of 5k abstracts that can be used for gathering evidence, SciFact-Open Wadden et al. (2022) expands this evidence retrieval corpus to 500K abstracts, presenting a more challenging retrieval problem. Also in the biomedical domain, COVID-Fact Saakyan et al. (2021) consists of over 4K claims extracted from Reddit. Lu et al. Lu et al. (2023) introduce SciTab, which requires verifying computer science claims centrally using tables extracted from papers. ClaimCheck Ou et al. (2025) uses reviews of rejected NeurIPS submissions from OpenReview to build a corpus of 154 claims that emphasize identifying the weaknesses in scholarly claims. In contrast, Matter-of-Fact builds a corpus of 8.4k materials science claims for feasibility assessment that are generated from the nuanced results found in the full text of source articles (rather than abstracts), and where negative claims focus on being scientifically infeasible rather than factually incorrect.
Claim Verification Models:
@@
Our framing of feasibility detection is as temporally-filtered claim verification with a knowledge cutoff. More broadly, recent approaches to claim verification typically involve two key steps: evidence retrieval and fact checking. For the retrieval step, augmenting LLMs with retrieved documents Izacard et al. (2022) or knowledge bases Baek et al. (2023); Hang et al. (2024) can be effective for improving fact verification performance of models. $\text{Re}^{2}\text{G}$ Glass et al. (2022) extends the retrieval step with a trained reranker to achieve better retrieval performance for fact checking. Rani et al. Rani et al. (2023) propose a form of query expansion that generates claim-related questions as queries to retrieve supporting documents. For fact checking, some methods make use of structured knowledge representations such as knowledge graphs Dammu et al. (2024) and first-order-logic Wang and Shu (2023) to organize evidence and verify facts. End-to-end systems combine the entire retrieval and verification pipeline, such as ARSJoint Zhang et al. (2021) and SciClaims Ortega and Gómez-Pérez (2025). In this work we demonstrate similar retrieval-backed systems (with temporal filtering) for feasibility assessment, while also providing formal approaches based on code generation.
Scientific Discovery and Feasibility Assessment:
@@
Automated scientific discovery is frequently divided into two subfields: problem-specific methods (like AlphaFold Jumper et al. (2021) for protein structure prediction), and problem-general methods that work across a variety of problem types. Examples of problem-specific systems in the materials science domain include GNoME Merchant et al. (2023), a graph neural network (GNN) based method that discovered over 2.2 million new stable crystal structures, and Schmidt et al. Schmidt et al. (2023)’s method for using crystal-graph neural networks together with high-quality data for accurate stability prediction. The latter work screened 1 billion materials, discovering 150k+ stable compounds, and identified extreme-property materials like superconductors. Similarly, Chen et al. Chen et al. (2024) combine machine learning models with traditional physics-based models to discover compounds to which can potentially serve as solid electrolytes. These problem-specific methods can be applied to feasibility assessment by predicting highly specific properties of unknown materials. Matter-of-Fact works to bridge the gap between problem-specific methods and problem-general methods by providing a large set of claims across 4 broad and high-impact areas of materials science, each of which is likely to benefit from a variety of problem-specific methods to arrive at accurate feasibility assessments. As we empirically demonstrate in our modeling results, because Matter-of-Fact nominally requires a large set of capacities to solve, it is challenging benchmark for measuring a general capacity to assess feasibility over broad subdomains.
@@ 3 Dataset
We assembled a corpus of recent publicly-available materials science domain articles by crawling Arxiv for all papers within the materials science and superconductivity topics submitted on or after January 2022, resulting in a total of 24k articles. Articles were then filtered based on specific inclusion criteria. First, articles that were not licensed using a specific permissive license (Creative Commons-By Attribution-4.0) were removed. Second, to prevent having to use a PDF-to-text conversion pipeline (which can have limited quality on complex tables, chemical formulas, mathematics, and other artifacts found within materials science articles), we further filtered to include only articles with  source available. Papers with long source ($>$30k tokens) were also removed (approximately 16% of articles). After initial filtering, 4.2k articles remained. Our focus in this work is specifically in four high-impact subdomains: superconductors, semi-conductors, batteries, and aerospace materials. To identify articles within these topics, we performed topic labeling of each abstract using gpt-4o-mini with a prompt that emphasized identifying articles within these 4 focus areas. We then sampled 500 total articles (125 from each topic) to use for claim generation.
Initial Claim Generation:
@@
Claims were generated by providing the full-text ( source) of each paper in a prompt, together with task instructions and JSON output format requirements. The model was instructed to generate matched pairs of claims – one true, and one that was clearly false or infeasible – and for each claim, to provide a list of supporting evidence, followed by an overall explanation as to why the evidence supports or refutes the claim. [^2] Scientific articles tend to express positive claims rather than negative claims. We follow the approach of Saakyan et al. Saakyan et al. (2021) to first extract positive claims, then automatically generate negative claims from these positive references. Claims were instructed to be stand-alone, and not make reference to the paper in the claim text (i.e. “Table 4 claims the boiling point of Material X is…”), so that they could be (in principle) solved without reference to the original source paper. Negative claims were instructed to be false or clearly infeasible (but not overly so), and not simply claims for which no evidence was available. Similarly, negative claims were instructed to use balanced language so as not to give away their true or false nature by particulars of wording, such as through use of negation markers (i.e. “Material X does not have…”), and to instead use neutral framings. In addition to the above constraints, claims were explicitly asked to be authored on two dimensions. The first asks for claims to specifically test either qualitative knowledge (e.g. “In Situation X, Phenonemon Y helps Material Z maintain its superconductivity”), or quantitative knowledge (e.g. “Material X superconducts at 77 Kelvin”). Second, claims were asked to be authored cross 4 main types: those that focus on experimental results, code/simulation results, theoretical results, or integrative methods across types.
Balanced Temporal Sets:
@@
Claims were temporally sorted into those from papers first appearing on Arxiv in 2022 (for training), those in 2023 (for validation), and the most recent claims from papers submitted between 2024 and April 2025 (for testing). For each set, we filtered claims such that equal numbers of true and false claims were present, to achieve a baseline (random chance) performance of 50%. Claims were also balanced such that equal numbers within the experimental, code, theory, and integrative categories appear within a given set. The final dataset includes a total of 8.4k claims, distributed as 1.4k claims for training, 2.5k for validation, and 4.4k for testing.
@@
Domain Expert Validation:
To measure the quality of the claim generation process, a domain expert with a graduate degree in materials science was given each claim and its source paper, and independently asked to determine the validity of the claim. This was a challenging task, because the claims span broad areas of materials science that would be unusual for any single individual to have expertise within. The domain expert performed this task for 100 claims from the test set, and initially agreed in 93% of cases (while noting a further 3% of claims appeared to not meet criteria, such as explicitly referencing the original source paper). They were then provided with the LLM-generated labels and explanations, and asked to resolve disagreements (either LLM errors, or human error), noting that nearly all errors were a result of the domain expert missing difficult-to-find evidence on their first attempt, and ultimately reaching 99% agreement after this resolution process. This empirically suggests that the overall data quality is high (96% after discounting data not meeting generation criteria). Before resolution, interrater agreement using Cohen’s Kappa Cohen (1960) was $\kappa=0.86$, or strong agreement McHugh (2012), suggesting the claims are highly objective.
@@ 4.1 Feasibility Assessment Models
The language model is provided with a prompt that includes the claim, and a request to think and/or plan before responding in the style of Chain-of-Thought Wei et al. (2022). We also include two variations of this model. The first includes a 20-shot in-context learning example Brown et al. (2020), using 20 claim problems (together with their supporting facts and explanations) drawn from the training set, including balanced numbers of true and false claims. The second includes adding a reflection step Madaan et al. (2023) where, after the initial generation, the model then reflects on its response, then provides a final answer and explanation for the reasoning behind that answer.
Retrieval Augmented Generation (RAG):
@@
Using the claim text as a query, the model first retrieves the top K matching full-text snippets from scholarly scientific articles using the SemanticScholar API Kinney et al. (2023), where each snippet generally takes the form of a span of text (approximately 500 words in length) from an article indexed by SemanticScholar that most closely matches the query. To prevent temporal contamination with oracle knowledge, full-text snippets are filtered such that papers authored after the source paper for a given claim are not included in the search. For example, if a claim was derived from a paper first published on Arxiv in March 2024, then only papers authored in February 2024 or before will be included in the snippet search. The top 20 matching full-text snippets (sorted by the provided relevance score) are included in the language model prompt, in a retrieval-augmented-generation paradigm Lewis et al. (2020). The prompt for this model also includes a 20-shot ICL example, and request for chain-of-thought reasoning.
Code Generation (CodeScientist):
@@
This model is performed in two stages. During the first stage, the model is provided with the claim text, and prompted to generate a code-based experiment or simulation in Python that would produce useful evidence in supporting or refuting the claim. The code is then executed, and the code and execution results are provided to a second prompt with a request to generate an answer for the feasibility task as well as a supporting explanation. For code execution, we use the experiment execution portion of CodeScientist Jansen et al. (2025), which allows executing arbitrary Python code in a virtual sandbox on Modal.com, and supports installing external supporting libraries through PIP. While this execution pipeline stores and saves output streams (e.g. stdout/stderr), the model explicitly prompted to save a log of its work, as well as a final list of results, which are then provided back to the model to help make its final decision. For tractability, we run CodeScientist in a limited form due to its high overall cost (initially reported as \$4 per experiment), which would be intractible for the size of our dataset (\$$\approx\$16k$ for 4k test claims). Instead of 25 debug iterations, we run CodeScientist for a single iteration (without reflection), and reduce the experiment time limit from 6 hours to 10 minutes (or 31 total CPU-days across all test claims). The model is made aware of these limitations in the code generation prompt, and encouraged to design appropriately-scoped experiments and output to support the decision process.
@@ 4.2 Claim Verification Models
RAG (Temporally Unrestricted):
The retrieval-augmented generation model described above, but without temporal restrictions. For a given claim, snippets from any scientific article may be retrieved, including (potentially) the source article of the claim, or those that cite the source article.
Oracle Baseline:
The language model is provided both with the claim, as well as the original source paper the claim was derived from (in the form of the paper’s original  source retrieved from Arxiv) in a retrieval-augmented-generation paradigm. This baseline measures how well a model can verify the claim when provided with a source scientific article that directly speaks to that claim’s validity/feasibility.
Oracle (Human Domain Expert):
The domain expert evaluation, as described in Section 3.

### v2

@@ Abstract
Contemporary approaches to assisted scientific discovery use language models to automatically generate large numbers of potential hypothesis to test, while also automatically generating code-based experiments to test those hypotheses. While hypotheses can be comparatively inexpensive to generate, automated experiments can be costly, particularly when run at scale (i.e. thousands of experiments). Developing the capacity to filter hypotheses based on their feasibility would allow discovery systems to run at scale, while increasing their likelihood of making significant discoveries. In this work we introduce Matter-of-Fact, a challenge dataset for determining the feasibility of hypotheses framed as claims, while operationalizing feasibility assessment as a temporally-filtered claim verification task using backtesting. Matter-of-Fact includes 8.4k claims extracted from scientific articles spanning four high-impact contemporary materials science topics, including superconductors, semiconductors, batteries, and aerospace materials, while including qualitative and quantitative claims from theoretical, experimental, and code/simulation results. We show that strong baselines that include retrieval augmented generation over scientific literature and code generation fail to exceed 72% performance on this task (chance performance is 50%), while domain-expert verification suggests nearly all are solvable – highlighting both the difficulty of this task for current models, and the potential to accelerate scientific discovery by making near-term progress. [^1] Benchmark, models, and claim extraction system:    https://github.com/cognitiveailab/matter-of-fact
@@ 1 Introduction
Generating data to test feasibility assessment is challenging, as (by nature) the experimental results of proposed hypotheses are as yet unknown, which makes gold labels for determining whether a hypothesis is feasible or infeasible effectively unavailable. To address this challenge, we operationalize feasibility assessment as a temporally-filtered claim verification task using the concept of backtesting. As with conventional claim verification tasks (e.g. Thorne et al., 2018), we generate a corpus of claims extracted from recent scientific literature – however, in addition, each claim is paired with a “knowledge cut-off date”, which is the date that the paper the claim was generated from was first authored. When performing the feasibility assessment task, models are allowed to use any information available before the source paper was authored to assess feasibility, essentially rewinding into the past to predict future results. In this way, models are provided with knowledge up to (for example) 2023, and must use that knowledge (through a combination of literature search, small-scale code-based experimentation, world modeling, or other methods) to predict whether genuine results (and artificially-generated infeasible results) from 2024 onward are feasible or infeasible.
@@
Matter-Of-Fact (This work) Mat. Sci 8.4k Paper full-text Extract positive, generate infeasible
@@ 2 Related Work
Scientific Claim Verification Datasets: The scientific claim verification task requires a model to determine whether a claim (typically extracted from a scientific paper) is true or false, either by leveraging its pretrained scientific knowledge or retrieving evidence from a corpus, with a selection of scientific claim verification benchmarks shown in Table 1. SciFact Wadden et al. (2020) contains 1.4K biomedical-domain claims generated by showing citances (sentences that cite a paper and describe its contribution) to human annotators, who were then asked to generate associated claims. Where SciFact pairs claims with a set of 5k abstracts that can be used for gathering evidence, SciFact-Open Wadden et al. (2022) expands this evidence retrieval corpus to 500K abstracts, presenting a more challenging retrieval problem. Also in the biomedical domain, COVID-Fact Saakyan et al. (2021) consists of over 4K claims extracted from Reddit. Lu et al. Lu et al. (2023) introduce SciTab, which requires verifying computer science claims centrally using tables extracted from papers. ClaimCheck Ou et al. (2025) uses reviews of rejected NeurIPS submissions from OpenReview to build a corpus of 154 claims that emphasize identifying the weaknesses in scholarly claims. In contrast, Matter-of-Fact builds a corpus of 8.4k materials science claims for feasibility assessment that are generated from the nuanced results found in the full text of source articles (rather than abstracts), and where negative claims focus on being scientifically infeasible rather than factually incorrect.
@@
Claim Verification Models: Our framing of feasibility detection is as temporally-filtered claim verification with a knowledge cutoff. More broadly, recent approaches to claim verification typically involve two key steps: evidence retrieval and fact checking. For the retrieval step, augmenting LLMs with retrieved documents Izacard et al. (2022) or knowledge bases Baek et al. (2023); Hang et al. (2024) can be effective for improving fact verification performance of models. $\text{Re}^{2}\text{G}$ Glass et al. (2022) extends the retrieval step with a trained reranker to achieve better retrieval performance for fact checking. Rani et al. Rani et al. (2023) propose a form of query expansion that generates claim-related questions as queries to retrieve supporting documents. For fact checking, some methods make use of structured knowledge representations such as knowledge graphs Dammu et al. (2024) and first-order-logic Wang and Shu (2023) to organize evidence and verify facts. End-to-end systems combine the entire retrieval and verification pipeline, such as ARSJoint Zhang et al. (2021) and SciClaims Ortega and Gómez-Pérez (2025). In this work we demonstrate similar retrieval-backed systems (with temporal filtering) for feasibility assessment, while also providing formal approaches based on code generation.
@@
Scientific Discovery and Feasibility Assessment: Automated scientific discovery is frequently divided into two subfields: problem-specific methods (like AlphaFold Jumper et al. (2021) for protein structure prediction), and problem-general methods that work across a variety of problem types. Examples of problem-specific systems in the materials science domain include GNoME Merchant et al. (2023), a graph neural network (GNN) based method that discovered over 2.2 million new stable crystal structures, and Schmidt et al. Schmidt et al. (2023)’s method for using crystal-graph neural networks together with high-quality data for accurate stability prediction. The latter work screened 1 billion materials, discovering 150k+ stable compounds, and identified extreme-property materials like superconductors. Similarly, Chen et al. Chen et al. (2024) combine machine learning models with traditional physics-based models to discover compounds to which can potentially serve as solid electrolytes. These problem-specific methods can be applied to feasibility assessment by predicting highly specific properties of unknown materials. Matter-of-Fact works to bridge the gap between problem-specific methods and problem-general methods by providing a large set of claims across 4 broad and high-impact areas of materials science, each of which is likely to benefit from a variety of problem-specific methods to arrive at accurate feasibility assessments. As we empirically demonstrate in our modeling results, because Matter-of-Fact nominally requires a large set of capacities to solve, it is challenging benchmark for measuring a general capacity to assess feasibility over broad subdomains.
@@ 3 Dataset
Inclusion Criteria: We assembled a corpus of recent publicly-available materials science domain articles by crawling Arxiv for all papers within the materials science and superconductivity topics submitted on or after January 2022, resulting in a total of 24k articles. Articles were then filtered based on specific inclusion criteria. First, articles that were not licensed using a specific permissive license (Creative Commons-By Attribution-4.0) were removed. Second, to prevent having to use a PDF-to-text conversion pipeline (which can have limited quality on complex tables, chemical formulas, mathematics, and other artifacts found within materials science articles), we further filtered to include only articles with  source available. Papers with long source (>>30k tokens) were also removed (approximately 16% of articles). After initial filtering, 4.2k articles remained. Our focus in this work is specifically in four high-impact subdomains: superconductors, semi-conductors, batteries, and aerospace materials. To identify articles within these topics, we performed topic labeling of each abstract using gpt-4o-mini with a prompt that emphasized identifying articles within these 4 focus areas. We then sampled 500 total articles (125 from each topic) to use for claim generation.
@@
Initial Claim Generation: Claims were generated by providing the full-text ( source) of each paper in a prompt, together with task instructions and JSON output format requirements. The model was instructed to generate matched pairs of claims – one true, and one that was clearly false or infeasible – and for each claim, to provide a list of supporting evidence, followed by an overall explanation as to why the evidence supports or refutes the claim. [^2] Scientific articles tend to express positive claims rather than negative claims. We follow the approach of Saakyan et al. Saakyan et al. (2021) to first extract positive claims, then automatically generate negative claims from these positive references.
Claims were instructed to be stand-alone, and not make reference to the paper in the claim text (i.e. “Table 4 claims the boiling point of Material X is…”), so that they could be (in principle) solved without reference to the original source paper. Negative claims were instructed to be false or clearly infeasible (but not overly so), and not simply claims for which no evidence was available. Similarly, negative claims were instructed to use balanced language so as not to give away their true or false nature by particulars of wording, such as through use of negation markers (i.e. “Material X does not have…”), and to instead use neutral framings. In addition to the above constraints, claims were explicitly asked to be authored on two dimensions. The first asks for claims to specifically test either qualitative knowledge (e.g. “In Situation X, Phenonemon Y helps Material Z maintain its superconductivity”), or quantitative knowledge (e.g. “Material X superconducts at 77 Kelvin”). Second, claims were asked to be authored cross 4 main types: those that focus on experimental results, code/simulation results, theoretical results, or integrative methods across types.
@@
Balanced Temporal Sets: Claims were temporally sorted into those from papers first appearing on Arxiv in 2022 (for training), those in 2023 (for validation), and the most recent claims from papers submitted between 2024 and April 2025 (for testing). For each set, we filtered claims such that equal numbers of true and false claims were present, to achieve a baseline (random chance) performance of 50%. Claims were also balanced such that equal numbers within the experimental, code, theory, and integrative categories appear within a given set. The final dataset includes a total of 8.4k claims, distributed as 1.4k claims for training, 2.5k for validation, and 4.4k for testing.
@@
Domain Expert Validation: To measure the quality of the claim generation process, a domain expert with a graduate degree in materials science was given each claim and its source paper, and independently asked to determine the validity of the claim. This was a challenging task, because the claims span broad areas of materials science that would be unusual for any single individual to have expertise within. The domain expert performed this task for 100 claims from the test set, and initially agreed in 93% of cases (while noting a further 3% of claims appeared to not meet criteria, such as explicitly referencing the original source paper). They were then provided with the LLM-generated labels and explanations, and asked to resolve disagreements (either LLM errors, or human error), noting that nearly all errors were a result of the domain expert missing difficult-to-find evidence on their first attempt, and ultimately reaching 99% agreement after this resolution process. This empirically suggests that the overall data quality is high (96% after discounting data not meeting generation criteria). Before resolution, interrater agreement using Cohen’s Kappa Cohen (1960) was $\kappa=0.86$, or strong agreement McHugh (2012), suggesting the claims are highly objective.
@@ 4.1 Feasibility Assessment Models
Chain-of-Thought (CoT), ICL, Reflection: The language model is provided with a prompt that includes the claim, and a request to think and/or plan before responding in the style of Chain-of-Thought Wei et al. (2022). We also include two variations of this model. The first includes a 20-shot in-context learning example Brown et al. (2020), using 20 claim problems (together with their supporting facts and explanations) drawn from the training set, including balanced numbers of true and false claims. The second includes adding a reflection step Madaan et al. (2023) where, after the initial generation, the model then reflects on its response, then provides a final answer and explanation for the reasoning behind that answer.
@@
Retrieval Augmented Generation (RAG): Using the claim text as a query, the model first retrieves the top K matching full-text snippets from scholarly scientific articles using the SemanticScholar API Kinney et al. (2023), where each snippet generally takes the form of a span of text (approximately 500 words in length) from an article indexed by SemanticScholar that most closely matches the query. To prevent temporal contamination with oracle knowledge, full-text snippets are filtered such that papers authored after the source paper for a given claim are not included in the search. For example, if a claim was derived from a paper first published on Arxiv in March 2024, then only papers authored in February 2024 or before will be included in the snippet search. The top 20 matching full-text snippets (sorted by the provided relevance score) are included in the language model prompt, in a retrieval-augmented-generation paradigm Lewis et al. (2020). The prompt for this model also includes a 20-shot ICL example, and request for chain-of-thought reasoning.
@@
Code Generation (CodeScientist): This model is performed in two stages. During the first stage, the model is provided with the claim text, and prompted to generate a code-based experiment or simulation in Python that would produce useful evidence in supporting or refuting the claim. The code is then executed, and the code and execution results are provided to a second prompt with a request to generate an answer for the feasibility task as well as a supporting explanation. For code execution, we use the experiment execution portion of CodeScientist Jansen et al. (2025), which allows executing arbitrary Python code in a virtual sandbox on Modal.com, and supports installing external supporting libraries through PIP. While this execution pipeline stores and saves output streams (e.g. stdout/stderr), the model explicitly prompted to save a log of its work, as well as a final list of results, which are then provided back to the model to help make its final decision. For tractability, we run CodeScientist in a highly limited form due to its high overall cost (initially reported as \$4 per experiment), which would be intractable for the size of our dataset ($\approx\mathdollar 16k$ for 4k test claims). Instead of 25 debug iterations, we run CodeScientist for a single iteration (without reflection), and reduce the experiment time limit from 6 hours to 10 minutes (or 31 total CPU-days across all test claims). The model is made aware of these limitations in the code generation prompt, and encouraged to design appropriately-scoped experiments and output to support the decision process.
@@ 4.2 Claim Verification Models
RAG (Temporally Unrestricted): The retrieval-augmented generation model described above, but without temporal restrictions. For a given claim, snippets from any scientific article may be retrieved, including (potentially) the source article of the claim, or those that cite the source article.
Oracle Baseline: The language model is provided both with the claim, as well as the original source paper the claim was derived from (in the form of the paper’s original  source retrieved from Arxiv) in a retrieval-augmented-generation paradigm. This baseline measures how well a model can verify the claim when provided with a source scientific article that directly speaks to that claim’s validity/feasibility.
Oracle (Human Domain Expert): The domain expert evaluation, as described in Section 3.